The perimeter of a rectangle is 56 inches. The ratio of the length to the width is 6:1. What is the length and the width.
Answer by mananth(16946)

ratio of length to width is 6:1
let the commoon multiple be x
length = 6x
width =x
Perimeter = 2(L+W)
2(6x+x) = 56
7x= 56/2
7x=28
/7
x= 4
so length = 6x= 6*4 24 in
Width = x = 4
2*4 + 24*2= 56